Continuous RVs + Conditioned. The PDF of a random variable X is given by: ( ) ?? + 0.4, 0 ≤ ? ≤ 2 0, ??ℎ?????? a. Find the value C that makes fX(x) a valid PDF. (Hint: Draw the PDF to start.) b. Find and sketch the CDF, FX(x). c. Calculate P[X > 1] d. Let A be the event that X > 1. Find and sketch the conditional PDF fX|A(x|A).
Answer #1

solving we get

C = 0.1

b)

cdf = P(X < x) =

c)

P(X > 1) = 1- P(X < 1)

= 1- 0.45

= 0.55

d)

P(X| X> 1) = (0.4 + 0.1x)/ 0.55